# Forecasting Melting Points in Svalbard, Norway Using Quantile Gradient Boosting and Adaptive Conformal Prediction Regions

### v1

@@ 6 Forecasts and Estimated Uncertainty
Figure 7: Mean forecasted temperature (horizontal axis) versus the percentage of cases in each of 20 equal-width bins where the true future temperature exceeded $0^{\circ}\mathrm{C}$ (vertical axis). Filled circles correspond to bins with mean forecasted temperature above $0^{\circ}\mathrm{C}$; open circles to those below. The solid blue line is a loess smoother. Constructed from the 2024 holdout dataset.
@@
Figure 7 provides a visual summary. The horizontal axis shows the mean of the daily forecasted temperatures for each of 20 equal-width bins. Because the actual temperatures are known for the 2024 holdout cases, the vertical axis shows the percentage of forecasts within each bin whose true future temperature exceeded 0. (The 14-day lag of each predictor ensures that these are future outcomes.)
The conclusions from Figure 7 are straightforward. False negatives are few—at most slightly under 20% of the forecasts in a bin—and occur mainly when the forecasted temperature lies between about $2^{\circ}\mathrm{C}$ and $1^{\circ}\mathrm{C}$. When the forecasted temperature exceeds $0^{\circ}\mathrm{C}$, more than 80% of the forecasts in a bin correspond to actual temperatures above 0, reaching 100% when the bin mean exceeds $4^{\circ}\mathrm{C}$. The large vertical gap between filled and open circles highlights the phase transition at $0^{\circ}\mathrm{C}$.
@@
Physical interpretation is subtle. When sea ice forms, its hexagonal lattice cannot readily incorporate salt ions. The excluded salt forms brine pockets and channels among the ice grains, leaving the ice nearly fresh water while the brine becomes increasingly saline. As freezing continues, the concentrated brine’s freezing point can fall well below $0^{\circ}\mathrm{C}$. Before that stage, a two-phase mixture of nearly pure ice and brine exists; melting reverses this process. Pure ice begins melting at $0^{\circ}\mathrm{C}$, whereas frozen brine melts at lower temperatures.
@@
The rates of melting and freezing depend on cloud cover, winds, and ocean currents that transport warmer or cooler water to local sites. Consequently, freezing and melting occur over a range of temperatures and time scales rather than at a fixed threshold. This interpretation aligns with Figure 7. From a policy perspective, even approximate information about when melting will occur can be valuable. Although figures like Figure 7 cannot be produced prospectively—since future temperatures are unknown—they can be constructed retrospectively using historical weather station data. Such analyses help stakeholders gauge likely rates of true positives and false negatives conditional on forecasted temperature. [^9] The plotted proportions should not be interpreted as probabilities. Temporal dependence persists among the observed temperatures within each bin, though its structure is not easily characterized. Observations are ordered by temperature rather than by time, and the time gaps among observations within a bin can range from days to months. For example, temperatures during late spring and early fall may be similar even though they occur several months apart.
Given this context, the role of conformal prediction regions can be revisited. For a new, unlabeled case T+1T\!+\!1, the true temperature will exceed the lower bound $L_{T+1}$ of its prediction region with probability at least $1-(\alpha/2)$. For $tau=.80$, that probability is .90. Moreover, because the conformal prediction regions are adaptive, they provide case-specific information: as shown earlier, forecasts above $0^{\circ}\mathrm{C}$ on the average yield narrower prediction regions than those for colder conditions.
@@ 7 Discussion
The analysis of 2 p.m. temperatures in Svalbard, Norway, combines three statistical traditions in an unusual way: supervised learning with a quantile as the estimation target, time-series analysis making use of credible holdout samples, and adaptive conformal prediction regions constructed with a different flavor of supervised learning.
The statistical procedures were informed by extensive knowledge of the research site. Equally important was guidance from climate science in formulating the algorithms and interpreting their output. The aim was to contribute both to scientific understanding and to climate-adaptation policy.
A two-week forecasting lead time before the onset of widespread surface melting is a very short fuse but would give Arctic community administrators critical time to prepare for infrastructure, safety, and logistical challenges (Streletskiy, Shiklomanov and Christiansen, 2019; U.S. Arctic Research Commission, 2003). Local governments could limit heavy-vehicle use on thaw-sensitive roads, runways, and around pipelines; pre-position maintenance materials; and inspect culverts, bridges, and drainage channels before damage occurs (Arctic Research Consortium of the United States (ARCUS), 2021). Water managers could adjust reservoir levels or activate temporary treatment measures to mitigate siltation and contamination from meltwater inflows.
In communities dependent on ice roads, snowmobile trails, or frozen river crossings, early warnings would permit orderly resupply and fuel delivery before surface travel becomes unsafe (Chen et al., 2025). Coastal and hillside settlements could prepare for increased risks of shoreline erosion or permafrost-related landslides (The Climate Institute and The Firelight Group, 2022). Health authorities might use the lead time to issue advisories on water quality or vector-borne disease risk as standing water accumulates. Because of Arctic warming, Iceland is now reported to host mosquitoes (Straker, 2025).
@@ 8 Conclusions
Appendix
Pseudocode 1: Constructing 2 p.m. Nonconformal Scores Input: For time $t=1,\ldots,T$, let $X_{t-14}$ denote the observed predictor values; $y_{t}$ denote the observed 2 p.m. solar-time temperature; $\tau$ the target quantile for the algorithmic fit; and $\alpha$ the value determining the desired coverage probability $1-\alpha$. 1. Using $\tau$, fit $y_{t}$ with $X_{t-14}$ via quantile gradient boosting, yielding algorithm $Q$. 2. With the trained boosting algorithm $\hat{Q}$ and $X_{t-14}$, compute fitted values $\hat{y}_{t}=\hat{Q}(X_{t-14})$. 3. Compute residuals $r_{t}=y_{t}-\hat{y}_{t}$. 4. If $r_{t}$ shows no temporal dependence, treat $r_{t}$ as exchangeable nonconformal scores. 5. If temporal dependence is present, fit a time-series procedure to $r_{t}$ to obtain white-noise residuals $\epsilon_{t}$. 6. Use the white-noise residuals $\epsilon_{t}$ as exchangeable nonconformal scores. 7. Using the upper and lower quantile bounds implied by $1-\alpha$, fit $\epsilon_{t}$ with $y_{t}$ and $X_{t-14}$ via quantile random forests $R$, saving the trained algorithm $\hat{R}$. Output: Save the nonconformal scores $\epsilon_{t}$, the trained quantile gradient-boosting algorithm $\hat{Q}$, and the trained quantile random-forests algorithm $\hat{R}$ for later use with a new unlabeled case.
Pseudocode 2: Forecasting 2 p.m. Temperatures Input: Let $X_{T+1}$ denote the 14-day-lagged predictor values for a new unlabeled case, and let $1-\alpha$ be the predetermined coverage probability. 1. Use the trained quantile gradient-boosting algorithm $\hat{Q}$ to obtain the forecast $\hat{y}_{T+1}=\hat{Q}(X_{T+1})$ for two weeks in the future. 2. Using the trained quantile random-forests algorithm $\hat{R}$ with predictors $\hat{y}_{T+1}$ and $X_{T+1}$, obtain the conformal upper and lower bounds for case $T+1$. 3. Extract from $\hat{R}$ the fitted values $q_{T+1,\alpha/2}$ and $q_{T+1,1-\alpha/2}$ as the lower and upper limits, respectively, of the prediction region at coverage level $1-\alpha$. Output: Construct the prediction interval: $\left[\,\hat{y}_{T+1}+q_{T+1,\alpha/2},\;\hat{y}_{T+1}+q_{T+1,1-\alpha/2}\,\right].$

### v2

@@ 6.1 Uncertainty Results
Figure 8: For the 2024 holdout sample, the mean forecasted temperature in a bin is on the horizontal axis. The percentage of cases in each of 20 equal-width bins having a true future temperature exceeding $0^{\circ}\mathrm{C}$ is on the vertical axis. Filled circles correspond to bins with mean forecasted temperature above $0^{\circ}\mathrm{C}$. Open circles correspond to those below. The solid blue line is a loess smoother.
@@
Figure 8 provides a visual summary of these ideas using the 2024 holdout test data. The horizontal axis shows the mean of the daily forecasted temperatures for each of 20 equal-width bins. Because the future temperatures are known for the 2024 holdout cases, the vertical axis shows the percentage of forecasts within each bin whose true future temperature exceeds $0^{\circ}\mathrm{C}$. (The 14-day lag of each predictor ensures that these are future outcomes.) The loess overlay is S-shaped, with the greatest rate of change between about $2^{\circ}\mathrm{C}$ and $2^{\circ}\mathrm{C}$.
The conclusions from Figure 8 are straightforward. The figure is a confirmation of the analyses above because there is strong evidence of forecasting skill in the holdout sample. As the forecasted temperature increases from about $10^{\circ}\mathrm{C}$ to about $10^{\circ}\mathrm{C}$, the percentage of cases above the true future melting temperature increases in a nearly monotonic manner from 0.0% to 100%.
Moreover, the percentages on the vertical axis in combination with the smoother confirm that true positives are far more common than false negatives. The largest percentage of false negatives is concentrated between about $2^{\circ}\mathrm{C}$ and $1^{\circ}\mathrm{C}$ where a little less than 20% of the future true temperature values exceed the melting point. As soon as the forecasted temperature exceeds $0^{\circ}\mathrm{C}$, more than 80% of the forecasts in a bin correspond to future temperatures above $0^{\circ}\mathrm{C}$, reaching 100% as the bin mean exceeds $4^{\circ}\mathrm{C}$. The large vertical gap between filled and open circles highlights the relatively abrupt phase transition centered on $0^{\circ}\mathrm{C}$ responsible for the dramatic increases in true positives. Because melting is such an important event, stakeholders might find Figure 8 encouraging. Some kinds of temperature forecasts might be accurate enough to be useful.
@@
But physical interpretations can be subtle. When sea ice forms, its hexagonal lattice cannot readily incorporate salt ions. The excluded salt forms brine pockets and channels among the ice grains, leaving the ice as nearly fresh water while the brine becomes increasingly salty. As freezing continues, the concentrated brine’s freezing point can fall well below $0^{\circ}\mathrm{C}$. Before that stage, a two-phase mixture of nearly pure ice and brine remains. Melting reverses this process: pure ice begins melting at $0^{\circ}\mathrm{C}$, whereas frozen brine melts at lower temperatures.
The rates of melting and freezing depend on cloud cover, winds, and ocean currents that transport warmer or cooler water to local sites. Consequently, freezing and melting occur over a range of temperatures and time scales rather than at a fixed threshold. This interpretation aligns with Figure 8.
@@
From a policy perspective, even approximate information about when melting will occur can be valuable. Although graphs like Figure 8 cannot be produced prospectively—since future temperatures are unknown—they can be constructed retrospectively using historical weather-station data. Such analyses help stakeholders gauge likely rates of true positives and false negatives conditional on forecasted temperature. For example, if a forecast exceeds $4^{\circ}\mathrm{C}$, the realized future temperature will almost certainly be $0^{\circ}\mathrm{C}$. [^13] The plotted percentages should not be interpreted as probabilities. Temporal dependence persists among the forecasted temperatures within each bin, although its structure is not easily characterized. The forecasts are ordered by temperature rather than by time, and the time gaps between forecasts within a bin can range from days to months. For example, forecasted temperatures during late spring and early fall may be similar even though they occur several months apart.
Given these results, the role of conformal prediction regions can be briefly revisited. For a new, unlabeled case T+1T{+}1, the true temperature will fall within its adaptive conformal prediction region with probability at least $1-\alpha$. When the lower bound of a prediction region is greater than zero, the probability that the future temperature will be above 0 is $1-\alpha/2$. For a specified coverage probability of 0.80, that probability would increase to 0.90. Coverage is gained because under some circumstances, the prediction region’s upper bound may not matter much. This approach could work well for Svalbard insofar as temperatures $0^{\circ}\mathrm{C}$ are far more concerning than lower temperatures. In the spirit of Figure 8, when the lower bound of the prediction region is greater than $0^{\circ}\mathrm{C}$, one may simply assume that the forecast of melting is correct.
@@ 7 Discussion
The analysis of 2 p.m. temperatures in Svalbard, Norway, combines three interlocking statistical traditions: gradient boosting with a quantile as the estimation target, an AR(1) model to construct exchangeable nonconformal scores, and adaptive conformal prediction regions computed with quantile random forests, which is robust to sparse data. To the best of the author’s knowledge, the combined use of three temporally separated samples, quantile gradient boosting, AR(1)-based whitening, and regime-specific adaptive conformal prediction has not previously appeared in the literature. The statistical procedures were informed by extensive knowledge of the research site. Equally important was guidance from climate science in the choice of algorithms and in interpreting their output. The aim was to contribute both to scientific understanding and to practical human adaptations to Arctic melting.
A two-week forecasting lead time before the onset of widespread surface melting is a very short fuse, but it would give Arctic community administrators critical time to prepare for infrastructure resilience, public safety, and logistical challenges (Streletskiy, Shiklomanov and Christiansen, 2019; U.S. Arctic Research Commission, 2003). Local governments could limit heavy-vehicle use on thaw-sensitive roads, runways, and around pipelines; pre-position maintenance materials; and inspect culverts, bridges, and drainage channels before damage occurs (Arctic Research Consortium of the United States (ARCUS), 2021). Water managers could adjust reservoir levels or activate temporary treatment measures to mitigate siltation and contamination from meltwater inflows.
In communities dependent on ice roads, snowmobile trails, or frozen river crossings, early warnings would permit orderly resupply and fuel delivery before surface travel becomes unsafe (Chen et al., 2025). Coastal and hillside settlements could prepare for increased risks of shoreline erosion or permafrost-related landslides (The Climate Institute and The Firelight Group, 2022). Health authorities might use the lead time to issue advisories on water quality or vector-borne disease risk as standing water accumulates. Because of Arctic warming, even Iceland is now reported to host mosquitoes (Straker, 2025).
@@ 8 Conclusions
Appendix A. Statistical Narrative
The adaptive conformal forecasting procedure proceeds sequentially in the following steps. Separate calibration and forecasting steps ensure that all prediction intervals are based solely on information available at the time of forecasting. The sequence below is written with R in mind but also applies to Python.
  1. 1.
    Data preparation. Subset the longitudinal Longyearbyen weather-station data into training data from 2023, calibration data from 2022, and test data from 2024. Specify a coverage probability $1-\alpha$. For each year, the response $y_{t}$ is the daily 2 p.m. air temperature in degrees Celsius. The predictors for each dataset $X_{t-14}$ are the same meteorological variables lagged by 14 days.
  2. 2.
    Algorithm training. With $\tau=0.60$, fit quantile gradient boosting to the 2023 training data (train), yielding a trained algorithm (gbm) and an optimal iteration count (best.iter).
  3. 3.
    Residuals for calibration. Using the 2022 calibration data (calibration), compute fitted values
    $\widehat{y}_{t}^{2022}=\texttt{predict}(\texttt{gbm},\texttt{calibration},\texttt{best.iter}),$
    and residuals
    $r_{t}^{2022}=y_{t}^{2022}-\widehat{y}_{t}^{2022}.$
  4. 4.
    Regime split. In response to subject-matter concerns, split the calibration residuals into warm and cool regimes using the fitted values:
    $\text{Warm if }\widehat{y}_{t}^{2022}>0,\qquad\text{Cool if }\widehat{y}_{t}^{2022}\leq 0.$
  5. 5.
    Temporal-dependence adjustment. Because of temporal dependence in $r_{t}^{2022}$, fit a first-order autoregressive model with intercept to the full residual sequence,
    $r_{t}^{2022}=\alpha+\phi r_{t-1}^{2022}+\varepsilon_{t}.$
    This produces a single set of AR(1) innovations. These innovations are then allocated to the warm and cool regimes according to the fitted-values split, which removes short-range serial dependence without altering the level or variability.
  6. 6.
    Innovations (nonconformal scores). Extract the AR(1) innovations $\varepsilon_{t}$ for each regime. These approximately white-noise innovations serve as the nonconformal scores.
  7. 7.
    Quantile-function estimation. For each regime, fit a quantile random forest (QRF) to the 2022 nonconformal scores, using the 2022 fitted values and the 2022 lagged predictors as covariates. QRF learns how the variability of the nonconformal scores depends on the fitted temperature level and on the lagged meteorological predictors.
  8. 8.
    Forecast generation. Using the 2024 test data (forecast), compute gbm forecasts
    $\widehat{y}_{t}^{2024}=\texttt{predict}(\texttt{gbm},\texttt{forecast},\texttt{best.iter}),$
    and assign each case to the warm or cool regime according to the sign of its forecast.
  9. 9.
    Nonconformal-score quantiles. For each 2024 case, supply its forecasted value and lagged predictors to the corresponding regime’s QRF to obtain predicted score quantiles $\widehat{q}_{\alpha/2}$ and $\widehat{q}_{1-\alpha/2}$.
  10. 10.
    Prediction intervals. Construct lower and upper adaptive conformal prediction bounds:
    $\mathrm{PI}_{\mathrm{lower}}=\widehat{y}_{t}^{2024}+\widehat{q}_{\alpha/2},\qquad\mathrm{PI}_{\mathrm{upper}}=\widehat{y}_{t}^{2024}+\widehat{q}_{1-\alpha/2}.$
    The interval widths reflect both the variability captured by QRF and any remaining dispersion in the nonconformal scores.
These linked procedures combine quantile gradient boosting, AR(1) whitening, and quantile-random-forest calibration to yield adaptive conformal prediction regions. The lower and upper bounds use the ranks of the nonconformal scores, while the magnitudes of those scores determine the lengths of the resulting prediction regions.
Exchangeability of Nonconformal Scores
The supervised learning algorithm is first fitted on the training data $\{(X_{t},Y_{t})\}_{t=1}^{T_{\mathrm{train}}}$. Predictions are then generated for the separate calibration data $\{X_{t}\}_{t=1}^{T}$, and the resulting residuals or their AR(1) innovations form the nonconformal scores $\{S_{t}\}_{t=1}^{T}$.
If $\{S_{t}\}_{t=1}^{T+1}$ are (approximately) exchangeable—perhaps after whitening— then, conditional on the fitted algorithm, the calibration scores and the new score $S_{T+1}$ are also exchangeable. Consequently,
$\Pr\!\left\{\,S_{T+1}\leq q_{1-\alpha}\big(\{S_{t}\}_{t=1}^{T}\cup\{S_{T+1}\}\big)\right\}\geq 1-\alpha,$
and the adaptive conformal prediction region attains marginal coverage of at least $1-\alpha$.
Exchangeability means that the joint distribution of the nonconformal scores is invariant to permutations of their indices. Although in practice the scores are realized in temporal order, their joint law would be unchanged had they been realized in any other order. That a new unlabeled case is realized at time T+1T{+}1 is immaterial. For the analyses of the Svalbard data, exchangeability depends on the whitening of the calibration data residuals. The residuals from the training data are not exchangeable, and neither are the residuals from the calibration data before whitening.